OpenStudy (anonymous):

If tan A = 18/17 , tan B = 1/35 , and 0°

hartnn (hartnn):

tan (A-B)=(tanA-tanB)/(1+tanA.tanB) does this help ?

OpenStudy (anonymous):

yes it sure does , i did with the same formula and got stuck in the middle !

hartnn (hartnn):

where exactly u stuck,because i am getting tan(A-B)=1

OpenStudy (anonymous):

@hartnn me too

OpenStudy (anonymous):

getting tan(A-B)=1

OpenStudy (anonymous):